# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v1

@@ II-C Standard RL and Maximum Entropy RL
In the reinforcement learning framework, the control system (1) is interpreted as a discrete-time deterministic environment, where $\mathcal{X}$ and $\mathcal{U}$ represent continuous state and action spaces, respectively. The application of control inputs to the system (1) corresponds to an agent interacting with this environment. Specifically, at each time step $t$, the agent observes the state $\mathbf{x}_{t}$ and executes an action $\mathbf{u}_{t}$ according to a policy $\pi$. Given a reward function $r:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$, which is typically designed to penalize deviations from the target state $\mathbf{x}_{g}$ or reference $\mathbf{x}_{t}^{\text{ref}}$ in accordance with Problem 1, the environment returns a scalar reward $r_{t}:=r(\mathbf{x}_{t},\mathbf{u}_{t})$ and transitions to the next state $\mathbf{x}_{t+1}$. Consequently, the interaction sequence is denoted by the transition tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\mathbf{x}_{t+1})$.
While classical control often employs deterministic laws, this article utilizes stochastic policies to facilitate exploration and robustness within the maximum entropy RL (MERL) framework. A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from a probability distribution. We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively. Unlike standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective:
@@
where $\gamma\in[0,1)$ is the discount factor, $\alpha>0$ is the temperature coefficient governing the trade-off between reward and exploration, and $\mathcal{H}(\pi(\cdot|\mathbf{x}))=\mathbb{E}_{\mathbf{u}\sim\pi(\cdot|\mathbf{x})}[-\log\pi(\mathbf{u}|\mathbf{x})]$ is the policy entropy.
The corresponding soft Q-function, $Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, represents the expected soft return, which comprises both rewards and entropy terms, starting from $(\mathbf{x}_{t},\mathbf{u}_{t})$ and following $\pi$ thereafter. The soft Bellman operator $\mathcal{T}^{\pi}$ is defined as:
@@
where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$. In the soft policy iteration (SPI) algorithm [19], the policy is optimized by alternating between soft policy evaluation, where $Q^{\pi}$ is updated via iterative application of $\mathcal{T}^{\pi}$, and soft policy improvement, which is equivalent to finding a policy $\pi$ that maximizes the soft Q-function augmented with an entropy term for $\mathbf{x}_{t}$:
@@
Although SPI converges to the optimal maximum entropy policy $\pi^{*}$ in tabular settings, practical control problems involve continuous spaces requiring function approximation. Thus, we parameterize the soft Q-function $Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})$ and the policy $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ using neural networks with parameters $\theta$ and $\phi$. We assume $\pi_{\phi}$ follows a diagonal Gaussian distribution with mean $\mu_{\phi}(\mathbf{x}_{t})$ and standard deviation $\sigma_{\phi}(\mathbf{x}_{t})$. Actions are sampled using the reparameterization trick: $\mathbf{u}_{t}=g_{\phi}(\mathbf{x}_{t};\epsilon_{t})$, where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m})$, ensuring that the objective remains differentiable with respect to $\phi$.
@@ III Learning Lyapunov Certificate
In this section, we develop the learning mechanism for the Lyapunov certificate within a model-free framework. We first introduce a sliding-window data collection strategy to satisfy multi-step trajectory requirements. Subsequently, we design a comprehensive loss function for the Lyapunov network, incorporating boundedness constraints and a multi-step stability loss. To enhance training efficiency and robustness, we employ trajectory-wise importance sampling for off-policy correction, a labeling mechanism for positive and negative samples, and a $\lambda$-weighted aggregation to balance the bias-variance trade-off.
@@ III-A Data Collection
Our algorithm is developed within the maximum entropy RL framework, where data is acquired through agent-environment interactions with the system (1). Unlike the standard SAC algorithm [19] that samples single-step transitions, we leverage the multi-step Lyapunov condition presented in (4), which necessitates data sequences spanning $n$ consecutive steps. We define a single-step interaction tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ represents the probability density of sampling action $\mathbf{u}_{t}$ at state $\mathbf{x}_{t}$. The collected multi-step data consists of sequences $\{\mathbf{d}_{t},\mathbf{d}_{t+1},\dots,\mathbf{d}_{t+n-1}\}$, which are stored in the experience replay buffer $\mathcal{D}$.
Data storage commences only after $n$ consecutive steps have been accumulated. Specifically, during the initial n1n-1 steps of any rollout, no data is committed to $\mathcal{D}$. Once the rollout length reaches or exceeds $n$, the collection process begins. As illustrated in Fig. 1, we employ a double-ended queue (deque) to temporarily buffer interaction data. During each agent-environment interaction, new data enters the deque. Once the deque reaches a length of $n$, the complete $n$-step sequence is transferred to $\mathcal{D}$, and the oldest data point is removed to maintain the deque size, enabling a continuous sliding window for data collection.
Figure 1: Schematic of the $n$-step sliding window data collection process. A double-ended queue (deque) of size $n$ is utilized to buffer consecutive interaction transitions $\mathbf{d}_{t}$, ensuring that complete $n$-step trajectories are available for the verification of the multi-step Lyapunov certificate.
@@ III-B Loss Function Design of Lyapunov Network
Similar to the soft Q-network and policy network, we introduce a Lyapunov network parameterized by $\psi$, denoted as $V_{\psi}$, to estimate the Lyapunov certificate. In designing the loss function for the Lyapunov certificate, we strictly adhere to the conditions established in Lemma 1, which encompass the boundedness and multi-step stability requirements.
(1) Boundedness Loss: Based on the first condition in Lemma 1, we construct the boundedness loss function to ensure the Lyapunov certificate is quadratically constrained:
@@
where $N$ denotes the number of multi-step trajectories in a batch, $i$ denotes the $i$-th multi-step sample in the batch, and $n$ represents the number of steps within each trajectory. This loss term explicitly constrains $V_{\psi}$ to satisfy the boundedness condition defined in (3).
(2) Multi-step Stability Loss: Consider a sampled multi-step trajectory $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ generated by a previous policy $\pi_{\phi_{\text{old}}}$. Since our goal is to train the Lyapunov certificate $V_{\psi}$ to be compatible with the current policy $\pi_{\phi_{\text{new}}}$ for subsequent policy updates, we must account for the off-policy nature of the data. We introduce trajectory-wise importance sampling to correct the distribution shift. For an $n$-step trajectory, the importance sampling ratio for the segment from $t$ to $t+k$ is defined as ($k=1,2,\dots,n-1$):
@@
Furthermore, we address the second condition in (3), namely the stability requirement. While (4) represents the ideal theoretical decay, empirical trajectories may not strictly satisfy the exponential stability bound in (5) during early stages of training. To bridge the gap between the stability condition and the stochastic samples, we introduce the concepts of positive and negative samples.
@@
Within an $n$-step sample, for a state $\mathbf{x}_{t+k}$, if it satisfies the exponential decay condition:
@@
Accordingly, we define the Exponential Stability Label (ESL) as:
@@
For trajectories composed of multi-step interaction data, we introduce a weighting factor $\lambda\in(0,1)$, inspired by the (truncated) $\lambda$-return in reinforcement learning [33], to balance the trade-off between bias and variance. Specifically, in the stability condition $V_{\psi}(\mathbf{x}_{t+k})\leq(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})$ from (4), the parameter $k$ determines the temporal horizon of the system’s evolution. When $k$ is small, the short trajectory horizon provides a limited view of the system’s long-term dynamics under policy $\pi_{\phi}$. Given that the Lyapunov network is randomly initialized and potentially inaccurate in the early stages of training, relying on short-horizon data can introduce significant estimation bias.
Conversely, a larger $k$ captures more extensive information regarding the trajectory evolution (extended horizon), thereby providing a stronger and more accurate learning signal that reduces bias. However, since each transition is sampled from the environment, every step introduces stochasticity; consequently, longer trajectories accumulate higher cumulative variance. Therefore, a small $k$ leads to high bias but low variance, while a large $k$ results in low bias but high variance. To optimize this trade-off, we employ the weighting factor $\lambda$ to aggregate multi-step information.
By integrating the positive/negative sample labels with the weighting factor $\lambda$, we construct the Lyapunov stability loss to enforce a rigorous correspondence between empirical samples and the stability condition (4):
@@
and the step-wise difference loss $\mathcal{L}_{\text{diff},k}^{i}(\psi)$ is defined as:
@@
(3) Combined Lyapunov Loss: By integrating the boundedness and multi-step stability losses, the total loss function for optimizing the Lyapunov network is formulated as:
@@ IV Learning Policy with Exponential Stability Constraints
This section presents the optimization framework for the proposed MSACL algorithm. We first establish the soft Q-network updates using multi-step Bellman residuals to enhance data efficiency. We then introduce a novel policy loss function that incorporates a stability advantage derived from the Lyapunov certificate, utilizing a PPO-style clipping mechanism to ensure robust convergence toward exponentially stable behavior. Finally, we describe the automated temperature adjustment and synthesize the complete actor-critic learning procedure.
@@ IV-A Loss Function Design of Soft Q Networks
Based on the soft Bellman operator (7), we define the sampled Bellman residual for the $i$-th transition within a multi-step sequence for the soft Q-network $Q_{\theta}$ as:
@@
where $Q_{\bar{\theta}}$ denotes the target network parameterized by $\bar{\theta}$, which is updated via polyak averaging to stabilize the learning process. Utilizing the collected multi-step sequences, the loss function for the soft Q-network is defined as the mean squared Bellman error (MSBE) across the entire trajectory:
@@
While standard implementations of maximum entropy RL algorithms, such as SAC [19], typically set the upper limit of the inner summation in (14) to zero (i.e., $n=1$) to utilize only single-step transitions, our formulation leverages the complete $n$-step sequence for parameter updates. Theoretically, minimizing the MSBE requires state-action pairs $(\mathbf{x}_{t},\mathbf{u}_{t})$ to be sampled from a distribution that ensures an unbiased stochastic gradient estimate. Specifically, the objective is given by:
$\mathcal{J}_{\text{SoftQ}}(\theta)=$ $\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\Big[Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})-\Big(r_{t}+\gamma\underset{\mathbf{x}_{t+1}\sim\rho_{\pi},\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}$
@@
Strictly speaking, the transitions within a multi-step sequence are temporally correlated, and utilizing the full sequence simultaneously deviates from the independent and identically distributed (i.i.d.) sampling requirement commonly assumed in stochastic optimization. However, this approach represents a deliberate trade-off between mathematical rigor and sample efficiency. By incorporating all transitions within each collected sequence into the loss function, we significantly enhance the data utilization rate and the information density of each training batch. Experimental evidence suggests that as long as the sequence length $n$ remains moderate, the bias introduced by these short-term correlations does not undermine the stability of the Q-function learning. In practice, the benefits of improved sample efficiency and accelerated convergence often outweigh the theoretical concerns regarding strict uniform sampling, ultimately contributing to more robust control performance.
@@
In the practical implementation of Algorithm 1, we utilize clipped double-Q learning with two networks, $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$, to mitigate the overestimation bias of soft Q-values, as motivated by the TD3 framework [25]. Furthermore, we incorporate the “delayed policy update” technique to ensure that the policy is updated against a more stable value function, thereby facilitating smoother convergence of the actor-critic process.
@@ IV-B Loss Function Design of Policy Network
This subsection details the optimization process for the policy network, which integrates the maximum entropy objective with a Lyapunov-based stability constraint. In the standard SAC framework [19], policy optimization aims to maximize the expected soft Q-value. For the $i$-th multi-step sequence, the SAC-component of the loss function is expressed as:
@@
To ensure the learned policy satisfies the exponential stability requirements, we utilize the learned Lyapunov certificate $V_{\psi}$ to guide the policy updates. Inspired by the Generalized Advantage Estimation (GAE) in PPO [24], we define a stability advantage function based on $V_{\psi}$, which serves as a metric for convergence performance:
@@
A positive stability advantage ($A_{t,k}\geq 0$) indicates that the system trajectory from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+k}$ strictly adheres to the exponential decay condition. Maximizing this advantage encourages the policy to select actions that accelerate the decay of $V_{\psi}$, thereby directly increasing the state convergence rate toward the equilibrium $\mathbf{x}_{g}$. Conversely, $A_{t,k}<0$ signifies a violation of the stability certificate, representing a disadvantage that the policy must learn to avoid. This stability-driven optimization paradigm aligns with recent developments in stable model-free control, such as D-learning [34, 35]. To robustly handle the bias-variance trade-off across varying horizons, we apply the weighting factor $\lambda$ to obtain the aggregated stability advantage:
@@
For a multi-step sample $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$, we define the policy importance sampling ratio based on the initial transition as $\rho^{i}(\phi)=\frac{\pi_{\phi}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}$. Incorporating the clipping mechanism to prevent excessively large policy updates, the final policy loss function is formulated as:
@@
It is important to distinguish between the two importance sampling (IS) strategies employed in this article. In (11), sequence-level IS (from $t$ to t+k1t+k-1) is used to train the Lyapunov certificate $V_{\psi}$. While sequence-level IS typically introduces higher variance, it provides the low-bias correction necessary for the Lyapunov network to accurately reflect the stability of the current policy $\pi_{\phi}$ using off-policy data. Since the policy parameters are fixed during the Lyapunov update phase, this variance remains manageable via the truncation in (10). However, when optimizing the policy $\pi_{\phi}$ itself, sequence-level IS would introduce prohibitive variance into the gradient estimates, often leading to training divergence. Consequently, following the design principle of PPO, we utilize only the first-step IS ratio $\rho(\phi)$ in (17). This approach effectively reduces the variance of the policy gradient, ensuring stable optimization while still benefiting from the multi-step stability information embedded within the weighted advantage $A_{t,\lambda}$.
IV-C Automated Temperature Adjustment
In maximum entropy RL, the temperature coefficient $\alpha$ in (6) is critical for balancing exploration and exploitation. Following the automated entropy adjustment mechanism introduced in [19], we update $\alpha$ by minimizing the following loss function:

### v2

@@ II-C Standard RL and Maximum Entropy RL
To begin the analysis, we interpret the control system (1) as a discrete-time environment with continuous state space $\mathcal{X}$ and action space $\mathcal{U}$ under the standard RL paradigm. At each time step $t$, the RL agent observes the state $\mathbf{x}_{t}\in\mathcal{X}$ and executes an action $\mathbf{u}_{t}\in\mathcal{U}$, causing the environment to transition to the next state $\mathbf{x}_{t+1}$ and yield a scalar reward $r_{t}=r(\mathbf{x}_{t},\mathbf{u}_{t})$. This reward is formulated to penalize state deviations from the target $\mathbf{x}_{g}$ (or reference signal $\mathbf{x}_{t}^{\text{ref}}$) while accounting for the associated control effort.
Furthermore, we employ stochastic policies in this article to facilitate exploration within the maximum entropy RL (MERL) framework. A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathcal{P}(\mathcal{U})$ denotes the set of probability distributions over $\mathcal{U}$, and $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from the probability distribution. We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively. In contrast to standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective [33, 34, 35]:
@@
where $\gamma\in[0,1)$ is the discount factor, $\alpha>0$ is the temperature coefficient, and $\mathcal{H}(\pi(\cdot|\mathbf{x}))=\mathbb{E}_{\mathbf{u}\sim\pi(\cdot|\mathbf{x})}[-\log\pi(\mathbf{u}|\mathbf{x})]$ is the policy entropy.
The corresponding soft Q-function, $Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, represents the expected soft return, which comprises both rewards and entropy terms. The soft Bellman operator $\mathcal{T}^{\pi}$ is defined as:
@@
where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$. In the soft policy iteration (SPI) algorithm [33], the policy is optimized by alternating between soft policy evaluation, where $Q^{\pi}$ is updated via $\mathcal{T}^{\pi}$, and soft policy improvement, which is equivalent to finding a policy $\pi$ that maximizes the soft Q-function for $\mathbf{x}_{t}$:
@@
Although SPI converges to the optimal policy $\pi^{*}$ in tabular settings, the control problems studied in this article involve continuous state and action spaces necessitating function approximation. Thus, we parameterize the soft Q-function $Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})$ and the policy $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ using neural networks with parameters $\theta$ and $\phi$. We assume $\pi_{\phi}$ follows a diagonal Gaussian distribution, and actions are sampled using the reparameterization trick [33]: $\mathbf{u}_{t}=g_{\phi}(\mathbf{x}_{t};\epsilon_{t})$, where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m})$, ensuring that the optimization objectives remain differentiable with respect to $\phi$.
@@ III Learning Lyapunov Certificate
This section details the model-free Lyapunov certificate learning mechanism. We first introduce a sliding-window strategy for multi-step data collection, and then design a loss function for the Lyapunov network (also referred to as the Lyapunov certificate), incorporating both boundedness and stability constraints.
@@ III-A Data Collection
In contrast to conventional Lyapunov-based RL algorithms [25, 36], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirement. Guided by the multi-step relationship (4), establishing exponential stability necessitates leveraging multi-step sequence samples from state trajectories. This sequential data structure provides a sufficient temporal horizon to validate the contraction properties of Lyapunov certificates.
To operationalize this, we define the fundamental transition tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$. Unlike standard replay buffers that store these tuples individually, we employ a double-ended queue (deque) as a sliding window to capture temporal correlations. Upon reaching a length of $n$, the complete multi-step sequence $\{\mathbf{d}_{t},\mathbf{d}_{t+1},\dots,\mathbf{d}_{t+n-1}\}$ is packaged and transferred to the replay buffer $\mathcal{D}$, while the oldest entry is concurrently removed to maintain the window size.
Figure 1: Schematic of $n$-step sliding window data collection. A double-ended queue (deque) of size $n$ buffers consecutive transitions $\mathbf{d}_{t}$ to provide complete multi-step trajectories.
@@ III-B Loss Function Design of Lyapunov Network
Similar to the soft Q-network and policy network, we introduce a Lyapunov network $V_{\psi}$ parameterized by $\psi$ to estimate the Lyapunov certificate. In designing the loss function, We strictly adhere to the boundedness and multi-step stability conditions in Lemma 1.
(1) Boundedness Loss. Based on the boundedness condition in (3), we construct the boundedness loss function to ensure the Lyapunov certificate to be quadratically constrained:
@@
where $N$ denotes the batch size of multi-step sequence samples, $i$ denotes the $i$-th sample, and $n$ is the sequence length. This loss enforces the boundedness condition on $V_{\psi}$.
(2) Multi-step Stability Loss. A major challenge in learning Lyapunov certificates from off-policy data is the discrepancy between the behavior policy $\pi_{\phi_{\text{old}}}$ and the target policy $\pi_{\phi}$. To mitigate this discrepancy, we employ a sequence-wise importance sampling strategy. Specifically, for a multi-step sequence $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ generated by $\pi_{\phi_{\text{old}}}$, we compute the cumulative importance sampling ratio for the transition from $t$ to $t+k$ ($k=1,\dots,n-1$) as:
@@
On the other hand, learning Lyapunov certificates directly from off-policy data may not be accurate, as empirical sequences in the early stages of training often fail to satisfy the exponential decay condition (5), leading to unreliable certificate estimation. To establish a rigorous correspondence between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism. Specifically, we introduce positive and negative samples to bridge the gap between theory and practice:
@@
Within an $n$-step sequence sample, for the state $\mathbf{x}_{t+k}$, if it satisfies the exponential decay condition:
@@
Accordingly, we define the Exponential Stability Label as:
@@
Furthermore, to balance the bias-variance trade-off in multi-step sequences, we introduce a weighting factor $\lambda\in(0,1)$, inspired by the (truncated) $\lambda$-return in classic RL literature [40]. Specifically, for the stability condition $V_{\psi}(\mathbf{x}_{t+k})\leq(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})$ from (4), small $k$ determines a short horizon with limited information. Given that the Lyapunov network is randomly initialized and potentially inaccurate in the early stages of training, relying on short-horizon data may introduce high estimation bias. Conversely, large $k$ captures extended information to reduce bias but accumulates environmental stochasticity into higher variance. To this end, we employ the weighting factor $\lambda$ to aggregate multi-step information and balance bias-variance trade-off.
Finally, by integrating the positive/negative sample labels with the weighting factor $\lambda$, we construct the Lyapunov stability loss to enforce a rigorous correspondence between empirical samples and the stability condition (4):
@@
where the step-wise difference loss $\mathcal{L}_{\text{diff},k}^{i}(\psi)$ is defined as:
@@
(3) Combined Lyapunov Loss. By integrating the boundedness loss and the multi-step stability loss, the comprehensive loss function for optimizing the Lyapunov network $V_{\psi}$ is formulated as:
@@ IV Learning Policy with Exponential Stability Constraints
This section details the MSACL optimization framework. We first establish soft Q-network updates using multi-step Bellman residuals to enhance data efficiency. Then we introduce a novel policy loss incorporating Lyapunov-based stability advantages with PPO-style clipping for robust exponential stabilization. Finally, we describe automated temperature adjustment and the complete actor-critic procedure.
@@ IV-A Loss Function Design of Soft Q Networks
Following the soft Bellman operator (7), we define the sampled Bellman residual for the $i$-th multi-step sequence sample of $Q_{\theta}$ as:
@@
where $Q_{\bar{\theta}}$ denotes the target network parameterized by $\bar{\theta}$. Utilizing collected multi-step sequence samples, the loss function for the soft Q-network is defined as the mean squared Bellman error (MSBE) across the entire sequence:
@@
In contrast to standard MERL algorithms such as SAC [33], which typically utilize single-step transition (i.e., $n=1$), our approach leverages the complete $n$-step sequence to minimize the following objective function:
$\mathcal{J}_{\text{SoftQ}}(\theta)=$ $\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\Big[Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})-\Big(r_{t}+\gamma\underset{\mathbf{x}_{t+1}\sim\rho_{\pi},\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}$ (15)
@@
Although minimizing (15) theoretically necessitates independent and identically distributed (i.i.d.) sampling for unbiased gradient estimation [41], our proposed multi-step approach strategically balances mathematical rigor with enhanced sample efficiency. Indeed, while $n$-step sequences introduce temporal correlations that deviate from strict i.i.d. assumptions, experimental evidence demonstrates that moderate sequence lengths maintain learning stability. Ultimately, the substantial gains in batch data information outweigh the minor bias introduced by short-term correlations, leading to accelerated convergence and more robust control performance.
@@ IV-B Loss Function Design of Policy Network
This subsection details the policy network optimization, which seamlessly integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
(1) Soft-Q Value Maximization. Following the policy optimization (8) in SPI, the first objective is to maximize the soft-Q value. For the $i$-th multi-step sequence sample $\{\mathbf{d}_{t}^{i},\dots,\mathbf{d}_{t+n-1}^{i}\}$, the SAC-style component of the loss function is expressed as:
@@
Note that $g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i})$ is obtained via the reparameterization trick [33].
(2) Stability-Aware Advantage Guidance. To ensure exponential stability, we attempt to utilize the learned Lyapunov certificate $V_{\psi}$ to guide policy updates. Inspired by the Generalized Advantage Estimation (GAE) in PPO [31], we transform the Lyapunov descent property into an active guidance signal. Specifically, we define the stability advantage function based on $V_{\psi}$ at $t+k$ as a metric for convergence performance:
@@
A positive stability advantage ($A_{t,k}\geq 0$) indicates that the system trajectory adheres to the exponential decay condition (4). Maximizing this advantage encourages actions that accelerate $V_{\psi}$ decay, thereby increasing the state convergence rate toward $\mathbf{x}_{g}$. Conversely, $A_{t,k}<0$ signifies a certificate condition violation that the policy must learn to avoid. This stability-aware optimization aligns with recent developments in stable model-free control, such as D-learning [26, 27, 28]. Similar to (11), in order to balance the bias-variance trade-off across varying horizons, we apply the weighting factor $\lambda$ to obtain the aggregated stability advantage:
@@
(3) Integrated Policy Optimization Objective. For the $i$-th multi-step sequence sample, we define the first-step importance sampling (IS) ratio as $\rho^{i}(\phi)=\frac{\pi_{\phi}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}$. By incorporating the aggregated stability advantage (17) with PPO-style clipping to prevent excessively large updates, the policy loss function is formulated as:
@@
Notably, we employ distinct importance sampling strategies for the critic and the actor. In (11), sequence-level IS provides low-bias correction for learning $V_{\psi}$ from off-policy data. This variance remains manageable via truncation in (10) since policy parameters are fixed during Lyapunov certificate updates. However, during policy optimization, where policy parameters are actively updated, sequence-level IS is avoided to prevent prohibitive variance and potential divergence. Following the PPO principle, we utilize only the first-step IS ratio $\rho^{i}(\phi)$ in (18). This ensures stable optimization while retaining multi-step stability information via the aggregated stability advantage $A_{t,\lambda}$.
IV-C Unified Actor-Critic Optimization Framework
Under the MERL paradigm, the temperature coefficient $\alpha$ in (6) is critical for balancing exploration and exploitation. Following the automated entropy adjustment mechanism introduced in [33], we update $\alpha$ by minimizing the following loss function: